Fermi National Laboratory
 

Operation Summary for December 17, 2010 - December 20, 2010

 

The Integrated Luminosity for the period from 12/13/10 to 12/20/10 was 66.31 Pb-1 (inverse picobarns). NuMI reported receiving 7.62E18 protons on target during this same period.

Friday Dec 17

The day shift began with the Tevatron (TEV) in a dry squeeze, with Operations monitoring the Recycler stash (322.5E10), with the Antiproton Source (Pbar) stacking (9.3E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-994 (JASMIN) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations put the TEV through a wet squeeze and then at 8:55 AM into HEP shot setup. Operations established store 8372 at 9:53 AM with an average initial luminosity (AIL) of 316.15E30. Operators reported at 11:50 AM that they had rebooted MECAR and brought a Booster RF station (BRF18) back online.

During the evening shift, Operations disabled the T-944 beam at 7:21 PM.

Saturday Dec 18

The midnight shift began with Operations monitoring store 8372 (74.55E30) and the stash (350.3E10), with Pbar stacking (13.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8372 at 1:11 AM. Operations reported a delay in shot setup due to trouble with a LRF3 switch tube rectifier card and control card. Operations established store 8374 at 6:05 AM with an AIL of 321.9E30. Operations established beam to T-944 at 6:12 AM.

During the day shift, Operations reported that the I- Source’s output was dropping (the H- Source was the source in use). Operators tuned on the I- Source through out the rest of the weekend to maintain output, according to instructions from Linac personnel.

During the evening shift, Operations disabled the T-944 beam at 7:54 PM. Operations terminated store 8374 at 9:21 PM, put the TEV into shot setup, and established store 8375 at 10:38 PM with an AIL of 327.3E30.

Sunday Dec 19

The midnight shift began with Operations monitoring store 8375 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-944 at 4:08 AM.

During the day shift, Operations terminated store 8375 at 1:42PM, put the TEV into shot setup, and established store 8376 at 3:08 PM with an AIL of 330.18E30.

During the evening shift, operators monitored the accelerator complex.

Monday Dec 20

The midnight shift began with Operations monitoring store 8376 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-944 at 4:07 PM. Operations terminated store 8376 at 5:58 AM, put the TEV into shot setup, and established store 8377 at 7:19 AM with an AIL of 350.6E30.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI and Booster personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/20/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory